OpenStudy (anonymous):

need help in algebra 2! im not good at math at all and absolutely dont understand algebra 2!

OpenStudy (anonymous):

write the equation of a polynomial that has zeros at -3 and 2.

Parth (parthkohli):

\[\textbf{HINT}\]If an equation has zeroes at, say, 4 and -3, it'd be:\[f(x) = (x - 4)(x -( 3)) = (x - 4)(x +3)\]

OpenStudy (anonymous):

2. n=4m^2-m+7m^4 a. n=4m^3+4m-2;cubic trinomial b. n=7m^4+4m^2-2;quartic trinomial c. n=4m^4 +8m^2-m; quartic trinomial d. n=3m^3+2m-5; cubic trinomial

Parth (parthkohli):

If you want a quadratic polynomial for what is needed, you'd have to take \(n\)'s degree two more than \(4m^2 - m + 7m^4\).

OpenStudy (anonymous):

3. (x^3+3x^2-x-3) divided by (x-1)

OpenStudy (callisto):

Start a new post for a new question.

OpenStudy (anonymous):

NO